Question 4 (ILOs: B1, B3, C3, D4)

Batelco company estimates that the demand for their products is

 

Q = 500 - 3P + 2Pr + 0.1Y

Where Q = quantity, Pr is the price of its rivals, and Y is income (currently, P = $10, Pr = $20, and Y = $6000)

 

a.    What is the price of elasticity of demand for Batelco?

 

 

 

 

 

 

b.    What is the income elasticity of demand for Batelco?

 

 

 

 

 

 

 

 

c.     What is the cross-price elasticity of demand between its products and its revival?

 

 

 

 

 

 

 

 

d.    What is the implicit assumption regarding the population in the market?  

1
Expert's answer
2021-10-25T09:19:33-0400

(a)

"Q=500-3P+2P_r+0.1Y"

"=500-3(10)+2(20)+0.1(6000)=1,110"

Price Elasticity of Demand"=b(\frac{P}{Q})"

"=-3(\frac{10}{1110})=0.03"


(b)

Income Elasticity of Demand:

"=d(\frac{Y}{Q})"

"=0.1(\frac{6000}{1110})"

"=0.54"


(c)

Cross price elasticity of demand:

"=c(\frac{P_r}{Q})"

"=2(\frac{20}{1110})=0.04"


(d)

Customers will buy from Batelco company because their products are technically superior.